# StreamFlow: Theory, Algorithm, and Implementation for High-Efficiency Rectified Flow Generation

### v1

@@ 1 Introduction
Rectified Flow (liu2022flow) and Flow Matching (lipman2022flow) are new generation methods that became popular about two years ago. Existing generation processes (such as diffusion (ho2020denoising; song2020denoising)) generally learn to gradually shift from a Gaussian noise distribution to the real distribution/image distribution. They usually require several dozen steps to predict the next noise from the previous noise. However, these two new methods attempt to provide a velocity field, allowing the model to directly predict the direction of the next distribution and the speed for each step, and then try to learn a more direct path between the distributions, enabling good results to be achieved in 4-step generation (yan2024perflow; liu2023instaflow).
@@
Previously, the generation based on diffusion models has had many limitations in terms of theory and mechanism, resulting in relatively slow generation efficiency. With the popularity of diffusion models in areas such as augmented/virtual reality, style transfer, and real-time action generation, there have been works called StreamDiffusion (kodaira2025streamdiffusion) that attempt to solve this problem. As shown in Fig. 1, they achieve this by creating a pipeline, distributing the asynchronous noise addition and noise removal processes to different pipelines, which makes it particularly fast in accelerating the generation of diffusion models.
@@
However, when we apply the related and similar acceleration theories and implementations to the flow model, many problems arise: (1) As mentioned in the first paragraph, their theories are not consistent. The content they generate at each step and the process of generating also differ. For example, they deal with time fields rather than noise. (2) In terms of scheduling strategies and other policies, they use the time step to dynamically divide the progress of the generated content and then guide the generation process. (3) When we attempt to make them compatible with the existing compilation-based acceleration methods (tensorrt2024), due to their distinct structures and the fact that they change the structure depending on the settings, this often leads to compilation failures.
@@
To address these issues, we first, based on the generation theory of the Flow model, attempted to create a batch processing of velocity fields. For different time steps as inputs, they were divided into different processing steps, conceptually treating them as a type of operation step. Secondly, we developed vectorized heterogeneous time-step batch processing for these operation steps. According to the sequence of these different times, they were dynamically filled into the pipeline for processing. Finally, we regarded the entire framework as a dynamic and structurally changing generation process, and further accelerated it using TensorRT (tensorrt2024) compilation.
@@
However, when we actually implemented it, some new challenges emerged [^1] These new challenges are specific to flow models and cannot be solved by any traditional acceleration framework.: (1) The time steps can be of varying lengths, unlike in the case of denoising where there is a minimum basic unit. This results in inconsistent and misaligned processing steps during the processing. (2) The existing scheduler wrapper is unable to handle the vectorized output and differentiable time steps. (3) Even if TensorRT cannot simply achieve the functions we envisioned, the compilation is not only slow but also prone to crashing.
@@ 2 Related Works
Flow Matching (FM) and Rectified Flow are recent paradigms that bridge diffusion models and continuous normalizing flows by directly regressing a velocity field $\mathbf{v}_{\theta}(\mathbf{x}_{t},t)$ that maps Gaussian noise $\mathbf{x}_{0}\sim\mathcal{N}(0,I)$ to data $\mathbf{x}_{1}\sim p_{\text{data}}$ via the ODE $d\mathbf{x}_{t}=\mathbf{v}_{\theta}(\mathbf{x}_{t},t)dt$ (lipman2022flow; albergo2022building). The key innovation lies in defining conditional probability paths that minimize transport cost. FM enables the use of optimal transport (straight-line) paths $\mathbf{x}_{t}=(1-t)\mathbf{x}_{0}+t\mathbf{x}_{1}$, trained by minimizing the regression objective $\mathcal{L}=\mathbb{E}_{t,\mathbf{x}_{0},\mathbf{x}_{1}}\|\mathbf{v}_{\theta}(\mathbf{x}_{t},t)-(\mathbf{x}_{1}-\mathbf{x}_{0})\|^{2}$ (lipman2022flow). This formulation allows for fast inference through simple Euler integration $\mathbf{x}_{1}\approx\mathbf{x}_{0}+\mathbf{v}_{\theta}(\mathbf{x}_{0},0)$ in as few as one to four steps. Rectified Flow extends this by introducing an iterative “reflow” procedure, denoted as $\mathbf{x}_{t}^{(k+1)}=(1-t)\mathbf{z}+t\cdot\text{ODE-Solve}(\mathbf{z},\mathbf{v}_{\theta}^{(k)})$, which progressively straightens the flow trajectories to reduce discretization error (liu2022flow). Building on this, methods like InstaFlow successfully apply text-conditioned reflow to large-scale Stable Diffusion models, achieving high-quality one-step generation (liu2023instaflow). These advancements lay the theoretical groundwork for our accelerated pipeline.
@@
Recent progress in generative modeling has been driven by diffusion, transformer, and flow-based architectures. DDPM (ho2020denoising) set the foundation for high-fidelity synthesis through iterative denoising, albeit with slow sampling. Latent Diffusion Models (LDM) (rombach2022high) reduce computation by operating in compressed latent space, while Diffusion Transformers (DiT) (peebles2023scalable) replace U-Nets with Vision Transformers to enhance scalability. Training these transformer-based models is further simplified by Representation Entanglement (REG) (wu2025representation), which speeds up convergence by injecting semantic tokens from foundation models. Beyond classical diffusion, hybrid designs such as VA-VAE (yao2025reconstruction) employ pretrained vision models to regularize high-dimensional latents, alleviating the reconstruction–generation trade-off. For fast sampling, MeanFlow (geng2025mean) enables one-step generation by estimating average trajectory velocity, and PeRFlow (yan2024perflow) accelerates inference via piecewise linear flow segments, functioning as a plug-and-play module. Additionally, Visual Autoregressive Modeling (VAR) (tian2024visual) emerges as a scalable alternative with a next-scale prediction scheme. These advances motivate our new rectified flow acceleration framework.
@@
To enable interactive generation (Fang_2025_ICCV; 11099440; fang2025signxfoundationmodelsign), researchers have explored inference acceleration via step reduction optimizations. Multi-step diffusion models can be distilled to few-step or one-step generation (liu2022flow; liu2023instaflow), while ODE-based solvers (liu2022flow; wu2025representation) reframe diffusion as neural ODEs for faster integration. Engineering optimizations like model quantization and TensorRT (tensorrt2024) can fuse layers and roughly double throughput by optimizing execution graphs. However, these assume fixed computation graphs and break under dynamic time steps or changing batch sizes common in RF-based methods.
@@ 3.1.1 Rectified Flow and Velocity Fields
Rectified Flow (liu2022flow; esser2024scalingrectifiedflowtransformers) formulates the generative process as learning a time-dependent velocity field $v_{\theta}(x_{t},t)$ that transports samples from a noise distribution $\pi_{0}$ to the data distribution $\pi_{1}$ along straight paths. The flow is defined by the ODE $\frac{dx_{t}}{dt}=v_{\theta}(x_{t},t)$ for $t\in[0,1]$. The model learns to predict the velocity field by minimizing:
@@ 3.1.2 Time-Windows Mechanism in PeRFlow
PeRFlow (yan2024perflow) introduces a time-windows mechanism that divides the time interval $[0,1]$ into $K$ non-overlapping windows $\mathcal{W}_{k}=[t_{k}^{s},t_{k}^{e}]$. For a given timestep $t_{c}\in\mathcal{W}_{k}$, the velocity is computed using window-specific parameters:
@@ 3.1.3 Challenges in Batch Acceleration
Prior work on accelerating diffusion models (lyu2022acceleratingdiffusionmodelsearly), such as StreamDiffusion (kodaira2025streamdiffusion), relies on batch denoising where all samples in a batch share the same timestep $t$. This homogeneous batching allows for $\mathbf{X}_{t^{\prime}}=f_{\theta}(\mathbf{X}_{t},t)$ where $\mathbf{X}_{t}\in\mathbb{R}^{B\times C\times H\times W}$. However, this approach fails for Rectified Flow models due to three fundamental challenges:
@@
Challenge 3: Dynamic Compilation. TensorRT and similar compilers assume static computational graphs. The time-window mechanism in Eq. LABEL:eq:time_windows creates dynamic structures that violate this assumption, causing compilation failures.
@@ 3.3.1 Decoupling Step Size from Velocity Field
The vanilla PeRFlow (yan2024perflow) scheduler processes timesteps sequentially, computing window parameters for each $t_{i}$ individually. We reformulate this as a vectorized operation that processes a batch of timesteps $\mathbf{t}=[t_{0},t_{1},\ldots,t_{B-1}]$ simultaneously.
@@ 3.4.2 Asynchronous Queue Processing
After the UNet forward pass, we apply the batched scheduler (Algorithm 1) to obtain updated latents $\mathcal{P}_{j}^{\prime}$. The pipeline buffer is then updated via a shift operation: $\mathcal{B}_{j+1}=[z_{new}^{\prime(S_{1})},z_{1}^{\prime(S_{2})},\ldots,z_{N-2}^{\prime(S_{N-1})}]$. The completed sample $z_{N-1}^{\prime(S_{N})}$ exits the pipeline for VAE (kingma2022autoencodingvariationalbayes) decoding. This asynchronous processing ensures continuous throughput: once the pipeline is full (after N1N-1 warmup iterations), each UNet call produces one complete sample.
@@ 3.5 Dynamic TensorRT Compilation
TensorRT (tensorrt2024) is a deep learning inference optimizer that compiles neural networks into optimized execution graphs. However, it assumes static inputs (chen2018tvmautomatedendtoendoptimizing): all samples in a batch must have identical shapes and processing paths. Our heterogeneous timestep batching violates this assumption because different timesteps may trigger different execution branches (shen2021nimbleefficientlycompilingdynamic) in the time-windows mechanism.
@@ 4 Experiments
Our work differs from previous acceleration frameworks which focus on optimizing different models(ma2023deepcacheacceleratingdiffusionmodels; li2024distrifusiondistributedparallelinference). Therefore, they do not compete with each other. Instead, our approach is compatible with the existing Rectified Flow models’ acceleration methods [^2] Accelerated methods for the traditional Rectified Flow model are generally non-transferable, require architecture changes, re-training or distillation.(liu2024instaflowstephighqualitydiffusionbased; yan2024perflowpiecewiserectifiedflow). They can be used in combination rather than being mutually exclusive. Therefore, we mainly focused on the effectiveness and reliability of our method, and conducted detailed ablation evaluations, power and memory efficiency evaluations, scalability robustness evaluations, and quality evaluations, etc. We test our pipeline on a workstation equipped with 8x NVIDIA Quadro RTX 8000 GPUs (46GB VRAM each), Intel Xeon Silver 4116 CPU @ 2.10GHz, and Ubuntu 24.04 LTS for image generation. We measured the inference time for an average of 100 images, following the convention of previous studies (kodaira2025streamdiffusion; feng2025streamdiffusionv2streamingdynamicinteractive).
@@
Baseline HF (von-platen-etal-2022-diffusers) 2.22 196.68 3696
Baseline Vanilla (yan2024perflow) 1.87 205.28 3788
@@ 4.1 Ablation Assessment
Our framework inherits two architectural-level optimizations from Xformer(zhang2024xformerhybridxshapedtransformer) and TensorRT (tensorrt2024). These can be switched, so there are two scenarios that need to be compared. Without any special indication, all are four-step inference speeds. VFB refers to Velocity Field Batching; taesd refers to the Tiny VAE (kingma2022autoencodingvariationalbayes) included in the acceleration framework; By default, all images are set to a size of 512*512.
As demonstrated in Figure 4, we conducted comprehensive ablation studies to validate each component’s contribution. The baseline performance shows that the vanilla implementation achieves 1.87 FPS while the Hugging Face pipeline reaches 2.22 FPS. When we incrementally add optimizations starting with Xformer architecture (Fig 4(a)) and taesd decoder (Fig 4(b)), we observe modest improvements to 2.3 FPS, confirming that tiny VAE decoding represents a significant bottleneck(rombach2022highresolutionimagesynthesislatent). The most substantial gains emerge when enabling VFB with taesd (Fig 4(c)), achieving 8.0 FPS—a 360% improvement that validates our core insight that batched velocity field (lipman2023flowmatchinggenerativemodeling) processing is essential for Rectified Flow models. Further enabling TensorRT compilation (Fig 4(d)) pushes performance to 10.05 FPS, while adding the compatibility layer (Fig 4(e)) introduces a minor trade-off (9.8 FPS) but enhances stability across diverse architectures. Our experiments reveal important scalability insights: even with 8 inference steps, our optimized pipeline (4.5 FPS, Fig 4(h)) outperforms the 4-step baseline (2.22 FPS), and the decoder batch size comparison (Fig 4(i)) shows that latent generation is no longer the bottleneck—the system is now limited by VAE decoding throughput.
@@ 4.2 Efficiency Assessment
Beyond raw speed improvements, we evaluate the resource efficiency of StreamFlow across power consumption and memory footprint metrics. Table 1 presents comprehensive measurements for our key configurations alongside baseline methods. The tensorrt+taesd+VFB setup achieves 10.05 FPS while consuming only 104-110W, compared to 205W for the vanilla baseline—representing a 47% reduction in power consumption while delivering 5.4× speedup, translating to approximately 11× better performance per watt (schwartz2019greenai). Even Xformer-based configurations demonstrate favorable power profiles, with xformer+taesd+VFB consuming 153W at 8.0 FPS versus the baseline’s 196W at 2.22 FPS. Memory efficiency constitutes another critical advantage of StreamFlow. Our optimized pipelines reduce peak memory usage from 3788MB (vanilla baseline) to 2666-3012MB across various configurations, achieving 20-30% memory savings through our batched velocity field processing strategy, which eliminates redundant intermediate tensor allocations. The TensorRT configurations achieve the lowest memory footprint (2666-2694MB) due to aggressive operator fusion and memory planning (chen2018tvmautomatedendtoendoptimizing; dao2022flashattentionfastmemoryefficientexact). Combined with throughput improvements, our framework enables serving 5-6× more requests per GPU, dramatically improving infrastructure utilization and cost efficiency in production environments.

### v2

@@ 1 Introduction
Rectified Flow (Liu et al., 2022) and Flow Matching (Lipman et al., 2022) are new generation methods that became popular about two years ago. Existing generation processes (such as diffusion (Ho et al., 2020; Song et al., 2020)) generally learn to gradually shift from a Gaussian noise distribution to the real distribution/image distribution. They usually require several dozen steps to predict the next noise from the previous noise. However, these two new methods attempt to provide a velocity field, allowing the model to directly predict the direction of the next distribution and the speed for each step, and then try to learn a more direct path between the distributions, enabling good results to be achieved in 4-step generation (Yan et al., 2024a; Liu et al., 2023).
@@
Previously, the generation based on diffusion models has had many limitations in terms of theory and mechanism, resulting in relatively slow generation efficiency. With the popularity of diffusion models in areas such as augmented/virtual reality, style transfer, and real-time action generation, there have been works called StreamDiffusion (Kodaira et al., 2025) that attempt to solve this problem. As shown in Fig. 1, they achieve this by creating a pipeline, by decoupling noise addition and denoising stages and mapping them to parallel pipeline queues, which makes it particularly fast in accelerating the generation of diffusion models.
@@
However, when attempting to transfer these diffusion-oriented acceleration frameworks to Rectified Flow models, we encounter fundamental mismatches at the level of representation, scheduling semantics, and execution structure: (1) As mentioned in the first paragraph, their theories are not consistent. The content they generate at each step and the process of generating also differ. For example, they deal with time fields rather than noise. (2) In terms of scheduling strategies and other policies, they use the time step to dynamically divide the progress of the content and then guide the generation process. (3) When we attempt to make them compatible with the existing compilation acceleration methods (NVIDIA Corporation, 2024), due to their distinct structures and the fact that they change the structure depending on the settings, this often leads to compilation failures.
@@
To address these issues, we first, based on the generation theory of the Flow model, attempted to create a batch processing of velocity fields. For different time steps as inputs, they were divided into different processing steps, conceptually treating them as a type of operation step. Secondly, we developed vectorized heterogeneous time-step batch processing for these operation steps. According to the sequence of these different times, they were dynamically filled into the pipeline for processing. Finally, we regarded the entire framework as a dynamic and structurally changing generation process, and further accelerated it using TensorRT (NVIDIA Corporation, 2024) compilation.
@@
However, when we actually implemented it, some new challenges emerged [^1] These new challenges are specific to flow models and cannot be solved by any traditional acceleration framework.:(1) Rectified Flow employs variable-length time steps, leading to misaligned and non-uniform processing stages. (2) The existing scheduler wrapper is unable to handle the vectorized output and differentiable time steps. (3) Even if TensorRT cannot simply achieve the functions we envisioned, the compilation is not only slow but also prone to crashing.
@@ 2 Related Works
Flow Matching (FM) and Rectified Flow are recent paradigms that bridge diffusion models and continuous normalizing flows by directly regressing a velocity field $\mathbf{v}_{\theta}(\mathbf{x}_{t},t)$ that maps Gaussian noise $\mathbf{x}_{0}\sim\mathcal{N}(0,I)$ to data $\mathbf{x}_{1}\sim p_{\text{data}}$ via the ODE $d\mathbf{x}_{t}=\mathbf{v}_{\theta}(\mathbf{x}_{t},t)dt$ (Lipman et al., 2022; Albergo and Vanden-Eijnden, 2022). The key innovation lies in defining conditional probability paths that minimize transport cost. FM enables the use of optimal transport (straight-line) paths $\mathbf{x}_{t}=(1-t)\mathbf{x}_{0}+t\mathbf{x}_{1}$, trained by minimizing the regression objective $\mathcal{L}=\mathbb{E}_{t,\mathbf{x}_{0},\mathbf{x}_{1}}\|\mathbf{v}_{\theta}(\mathbf{x}_{t},t)-(\mathbf{x}_{1}-\mathbf{x}_{0})\|^{2}$ (Lipman et al., 2022). This formulation allows for fast inference through simple Euler integration $\mathbf{x}_{1}\approx\mathbf{x}_{0}+\mathbf{v}_{\theta}(\mathbf{x}_{0},0)$ in as few as one to four steps. Rectified Flow extends this by introducing an iterative “reflow” procedure, denoted as $\mathbf{x}_{t}^{(k+1)}=(1-t)\mathbf{z}+t\cdot\text{ODE-Solve}(\mathbf{z},\mathbf{v}_{\theta}^{(k)})$, which progressively straightens the flow trajectories to reduce discretization error (Liu et al., 2022). Building on this, methods like InstaFlow successfully apply text-conditioned reflow to large-scale Stable Diffusion models, achieving high-quality one-step generation (Liu et al., 2023). These advancements lay the theoretical groundwork for our accelerated pipeline.
@@
Recent progress in generative modeling has been driven by diffusion, transformer, and flow-based architectures. DDPM (Ho et al., 2020) set the foundation for high-fidelity synthesis through iterative denoising, albeit with slow sampling. Latent Diffusion Models (LDM) (Rombach et al., 2022a) reduce computation by operating in compressed latent space, while Diffusion Transformers (DiT) (Peebles and Xie, 2023) replace U-Nets with Vision Transformers to enhance scalability. Training these transformer-based models is further simplified by Representation Entanglement (REG) (Wu et al., 2025), which speeds up convergence by injecting semantic tokens from foundation models. Beyond classical diffusion, hybrid designs such as VA-VAE (Yao et al., 2025) employ pretrained vision models to regularize high-dimensional latents, alleviating the reconstruction–generation trade-off. For fast sampling, MeanFlow (Geng et al., 2025) enables one-step generation by estimating average trajectory velocity, and PeRFlow (Yan et al., 2024a) accelerates inference via piecewise linear flow segments, functioning as a plug-and-play module. Additionally, Visual Autoregressive Modeling (VAR) (Tian et al., 2024) emerges as a scalable alternative with a next-scale prediction scheme. These advances motivate our new rectified flow acceleration framework.
@@
To enable interactive generation (Fang et al., 2025, 2026), researchers have explored inference acceleration via step reduction optimizations. Multi-step diffusion models can be distilled to few-step or one-step generation (Liu et al., 2022, 2023), while ODE-based solvers (Liu et al., 2022; Wu et al., 2025) reframe diffusion as neural ODEs for faster integration. Engineering optimizations like model quantization and TensorRT (NVIDIA Corporation, 2024) can fuse layers and roughly double throughput by optimizing execution graphs. However, these assume fixed computation graphs and break under dynamic time steps or changing batch sizes common in Flow methods.
@@ 3.1.1 Rectified Flow and Velocity Fields
Rectified Flow (Liu et al., 2022; Esser et al., 2024) formulates the generative process as learning a time-dependent velocity field $v_{\theta}(x_{t},t)$ that transports samples from a noise distribution $\pi_{0}$ to the data distribution $\pi_{1}$ along straight paths. The flow is defined by the ODE $\frac{dx_{t}}{dt}=v_{\theta}(x_{t},t)$ for $t\in[0,1]$. The model learns to predict the velocity field by minimizing:
@@ 3.1.2 Time-Windows Mechanism in PeRFlow
PeRFlow (Yan et al., 2024a) introduces a time-windows mechanism that divides the time interval $[0,1]$ into $K$ non-overlapping windows $\mathcal{W}_{k}=[t_{k}^{s},t_{k}^{e}]$. For a given timestep $t_{c}\in\mathcal{W}_{k}$, the velocity is computed using window-specific parameters:
@@ 3.1.3 Challenges in Batch Acceleration
Prior work on accelerating diffusion models (Lyu et al., 2022), such as StreamDiffusion (Kodaira et al., 2025), relies on batch denoising where all samples in a batch share the same timestep $t$. This homogeneous batching allows for $\mathbf{X}_{t^{\prime}}=f_{\theta}(\mathbf{X}_{t},t)$ where $\mathbf{X}_{t}\in\mathbb{R}^{B\times C\times H\times W}$. However, this approach fails for Rectified Flow models due to three fundamental challenges:
@@
Challenge 3: Dynamic Compilation. TensorRT and similar compilers assume static computational graphs. The time-window mechanism in Eq. 4 creates dynamic structures that violate this assumption, causing compilation failures.
@@ 3.3.1 Decoupling Step Size from Velocity Field
The vanilla PeRFlow (Yan et al., 2024a) scheduler processes timesteps sequentially, computing window parameters for each $t_{i}$ individually. We reformulate this as a vectorized operation that processes a batch of timesteps $\mathbf{t}=[t_{0},t_{1},\ldots,t_{B-1}]$ simultaneously.
@@ 3.4.2 Asynchronous Queue Processing
After the UNet forward pass, we apply the batched scheduler (Algorithm 1) to obtain updated latents $\mathcal{P}_{j}^{\prime}$. The pipeline buffer is then updated via a shift operation: $\mathcal{B}_{j+1}=[z_{new}^{\prime(S_{1})},z_{1}^{\prime(S_{2})},\ldots,z_{N-2}^{\prime(S_{N-1})}]$. The completed sample $z_{N-1}^{\prime(S_{N})}$ exits the pipeline for VAE (Kingma and Welling, 2022) decoding. This asynchronous processing ensures continuous throughput: once the pipeline is full (after N1N-1 warmup iterations), each UNet call produces one complete sample.
@@ 3.5 Dynamic TensorRT Compilation
TensorRT (NVIDIA Corporation, 2024) is a deep learning inference optimizer that compiles neural networks into optimized execution graphs. However, it assumes static inputs (Chen et al., 2018): all samples in a batch must have identical shapes and processing paths. Our heterogeneous timestep batching violates this assumption because different timesteps may trigger different execution branches (Shen et al., 2021) in the time-windows mechanism.
@@ 4 Experiments
Our work differs from previous acceleration frameworks which focus on optimizing different models(Ma et al., 2023; Li et al., 2024). Therefore, they do not compete with each other. Instead, our approach is compatible with the existing Rectified Flow models’ acceleration methods [^2] Accelerated methods for the traditional Rectified Flow model are generally non-transferable, require architecture changes, re-training or distillation.(Liu et al., 2024; Yan et al., 2024b). They can be used in combination rather than being mutually exclusive. Therefore, we mainly focused on the effectiveness and reliability of our method, and conducted detailed ablation evaluations, power and memory efficiency evaluations, scalability robustness evaluations, and quality evaluations, etc. We test our pipeline on a workstation equipped with 8x NVIDIA Quadro RTX 8000 GPUs (46GB VRAM each), Intel Xeon Silver 4116 CPU @ 2.10GHz, and Ubuntu 24.04 LTS for image generation. We measured the inference time for an average of 100 images, following the convention of previous studies (Kodaira et al., 2025; Feng et al., 2025).
@@
Baseline HF (von Platen et al., 2022) 2.22 196.68 3696
Baseline Vanilla (Yan et al., 2024a) 1.87 205.28 3788
@@ 4.1 Ablation Assessment
Our framework inherits two architectural-level optimizations from Xformer(Zhang et al., 2024) and TensorRT (NVIDIA Corporation, 2024). These can be switched, so there are two scenarios that need to be compared. Without any special indication, all are four-step inference speeds. VFB refers to Velocity Field Batching; taesd refers to the Tiny VAE (Kingma and Welling, 2022) included in the acceleration framework; By default, all images are set to a size of 512*512.
As demonstrated in Figure 4, we conducted comprehensive ablation studies to validate each component’s contribution. The baseline performance shows that the vanilla implementation achieves 1.87 FPS while the Hugging Face pipeline reaches 2.22 FPS. When we incrementally add optimizations starting with Xformer architecture (Fig 4(a)) and taesd decoder (Fig 4(b)), we observe modest improvements to 2.3 FPS, confirming that tiny VAE decoding represents a significant bottleneck(Rombach et al., 2022b). The most substantial gains emerge when enabling VFB with taesd (Fig 4(c)), achieving 8.0 FPS—a 360% improvement that validates our core insight that batched velocity field (Lipman et al., 2023) processing is essential for Rectified Flow models. Further enabling TensorRT compilation (Fig 4(d)) pushes performance to 10.05 FPS, while adding the compatibility layer (Fig 4(e)) introduces a minor trade-off (9.8 FPS) but enhances stability across diverse architectures. Our experiments reveal important scalability insights: even with 8 inference steps, our optimized pipeline (4.5 FPS, Fig 4(h)) outperforms the 4-step baseline (2.22 FPS), and the decoder batch size comparison (Fig 4(i)) shows that latent generation is no longer the bottleneck—the system is now limited by VAE decoding throughput.
@@ 4.2 Efficiency Assessment
Beyond raw speed improvements, we evaluate the resource efficiency of StreamFlow across power consumption and memory footprint metrics. Table 1 presents comprehensive measurements for our key configurations alongside baseline methods. The tensorrt+taesd+VFB setup achieves 10.05 FPS while consuming only 104-110W, compared to 205W for the vanilla baseline—representing a 47% reduction in power consumption while delivering 5.4× speedup, translating to approximately 11× better performance per watt (Schwartz et al., 2019). Even Xformer-based configurations demonstrate favorable power profiles, with xformer+taesd+VFB consuming 153W at 8.0 FPS versus the baseline’s 196W at 2.22 FPS. Memory efficiency constitutes another critical advantage of StreamFlow. Our optimized pipelines reduce peak memory usage from 3788MB (vanilla baseline) to 2666-3012MB across various configurations, achieving 20-30% memory savings through our batched velocity field processing strategy, which eliminates redundant intermediate tensor allocations. The TensorRT configurations achieve the lowest memory footprint (2666-2694MB) due to aggressive operator fusion and memory planning (Chen et al., 2018; Dao et al., 2022). Combined with throughput improvements, our framework enables serving 5-6× more requests per GPU, dramatically improving infrastructure utilization and cost efficiency in production environments.